In a right triangle, one acute angle is 54° more than twice the other. Find each acute angle.
Answer by ikleyn(52778)

.
Let x be that acute angle.  Then the condition says


x = 2*(90-x) + 54 degree,

x = 180 - 2x + 54

3x = 180 + 54

x = 60 + 18 = 78.


Answer.    78 degrees  and  90-78 = 12 degrees.